# Structured Agent Distillation for Large Language Model

### v2

@@ 3.5 Curriculum Sampling in Structured Distillation
$C(\tau)=\alpha\cdot\text{len}(r_{1:k})+\beta\cdot\text{len}(a_{1:m})+\gamma\cdot\text{entropy}(\pi_{T}(\tau)),$ (7)
where $\text{len}(r_{1:k})$ and $\text{len}(a_{1:m})$ denote the lengths of the reasoning and action segments, respectively, and $\text{entropy}(\pi_{T}(\tau))$ reflects teacher uncertainty. The weights $\alpha,\beta,\gamma$ balance their relative contributions. During training, trajectories are sorted by $C(\tau)$, allowing the model to start with simpler examples and gradually progress to more complex ones.
3.6 Training Algorithm
Algorithm 1 outlines our structured agent distillation process. The student policy $\pi_{\theta}$ learns to imitate the teacher $\pi_{T}$ across reasoning stages, guided by a curriculum scheduler $\mathcal{C}$ that samples increasingly complex trajectories. Each trajectory is tokenized into reasoning and action spans, and the student predicts tokens autoregressively. The objective aggregates reasoning, action imitation losses, and gradients are backpropagated to update $\theta$.
Algorithm 1 Structured Agent Distillation
@@ 3.6 Training Algorithm
2:for epoch = 1 to E do
3:  Sample trajectory $\tau\sim\mathcal{C}$
4:  Segment into $\tau^{(r)},\tau^{(a)}$ and tokenize $x_{1:T}$
5:  Compute segment masks $m_{r},m_{a}$
6:  Perform forward pass $\hat{x}=\pi_{\theta}(x)$
@@
8:  Backpropagate $\mathcal{L}_{\text{total}}$ and Update $\theta\leftarrow\theta-\nabla\mathcal{L}_{\text{total}}$
9:end for
4 Experiments
4.1 Experimental Setups
Agent Environments. We evaluate our framework on three agent environments: (i) ALFWorld ALFWorld20 for embodied instruction following, (ii) WebShop yao2022webshop for web-based tool planning, and (iii) HotPotQA-ReAct yao2023react for multi-hop QA with reasoning traces. Task-specific details are in Appendix B.
Models. We apply Structured Agent Distillation to compress ReAct-style GPT-2-1.5B teacher trajectories into student models of 120M, 340M, and 760M (GPT-2); OPT-13B into OPT students (1.3B, 2.7B, 6.7B); and LLaMA-13B into LLaMA-7B.
Baselines. Our goal is not to exhaustively benchmark all existing distillation methods, but rather to evaluate how span-level alignment improves student agent fidelity under comparable conditions. We compare against token-level baselines, primarily using MiniLLM gu2024minillm —a state-of-the-art student model distilled from GPT-2 via token-level imitation, which minimizes the KL divergence between student and teacher outputs at every decoding step.
4.2 Experimental Results and Analysis
@@
Table 3: Unified comparison of task success (↑), reasoning length (↓), CoT match (↑), and latency (↓) across ALFWorld, WebShop, and HotPotQA. The teacher is a GPT-2-1.5B ReAct-style agent. Students trained via Structured Agent Distillation consistently outperform token-level baselines gu2024minillm .
@@
Table 4: Unified comparison of task success (↑), reasoning length (↓), CoT match (↑), and latency (↓) across ALFWorld, WebShop, and HotPotQA. The teachers are OPT-13B and LLaMA-13B ReAct-style agents. Students trained via Structured Agent Distillation consistently outperform token-level baseliness gu2024minillm .
Model Task Success ↑ Reasoning Length ↓ CoT Match ↑ Latency ↓
@@
Toke-LLaMA-7B 64.2 59.3 71.5 41.1 37.5 43.2 73.0 68.3 81.2 6.7 6.1 5.2
@@
CoT Consistency. As shown in Table 3, our method achieves higher CoT match rates across all settings, demonstrating stronger structural alignment with teacher reasoning.
@@
Segment Mask Validation. Figure 7 (Appendix F.2) confirms that token-level masks align accurately with reasoning/action spans across environments.
@@
5 Scaling Analysis
Figure 2: Scaling behavior of student agents across model sizes. Top-Left: Task Success Rate (%). Top-Right: Reasoning Length (tokens). Bottom-Left: Chain-of-Thought (CoT) Match Rate (%). Bottom-Right: Latency (steps per episode). Structured Agent Distillation consistently outperforms MiniLLM-OPT baselines and better approaches teacher performance as model capacity increases.
We evaluate how Structured Agent Distillation scales by transferring GPT-2-1.5B teacher trajectories into student models with 120M, 340M, and 760M parameters, compared against corresponding MiniLLM-OPT-GPT2 baselines.
@@ 5 Scaling Analysis
Summary. Structured Agent Distillation scales effectively with student capacity, enhancing task success, planning efficiency, and structural reasoning alignment. Improvements are most pronounced at smaller scales (e.g., 120M, 340M), where it mitigates the performance degradation commonly observed with token-level imitation. Additional scaling results for OPT and LLaMA models are presented in Appendix G.
6 Ablation Studies
We conduct ablation studies to understand the contribution of each component in our Structured Agent Distillation framework. Specifically, we analyze the roles of reasoning supervision, action supervision, and span segmentation.
Table 5: Ablation study on ALFWorld using a 340M-parameter student model. Each variant controls whether reasoning, action supervision, and explicit span segmentation are enabled. We report task success rate, CoT match rate, and episode latency(steps).
Method Reason Action Segm Succ (%) CoT (%) Steps
@@ 6 Ablation Studies
Removing reasoning supervision ($\mathcal{L}_{\text{CoT}}$) significantly degrades CoT match and increases latency, indicating reduced planning coherence.
@@
These results demonstrate that our method goes beyond masked token-level imitation—segment-aware supervision captures structurally distinct signals crucial for reasoning-action alignment.
7 Discussion
@@
Loss Design and Supervision. Although our total loss aggregates $\mathcal{L}_{\text{CoT}}$ and $\mathcal{L}_{\text{Act}}$, they are not functionally redundant. As shown in ablation studies, removing either term leads to substantial drops in performance across CoT alignment, task success, and planning efficiency. This confirms that span-level supervision provides complementary guidance that token-level KL losses alone cannot capture.
@@
Limitations. Our current framework is designed for text-only agents operating under ReAct-style assumptions. It does not yet model multimodal spans (e.g., vision-language), nor does it account for agent memory or external feedback loops.
Future Work. We aim to extend span-aware supervision to multimodel agents, and explore alignment for more complex structures, such as hierarchical subgoals or latent plans. We plan to use diffusion models meng2024instructgie to generate small samples liu2023interpretable ; liu2021explainable ; liu2022efficient for multimodal agent research in healthcare chinta2025ai ; liu2024brain ; chinta2023optimization ; wang2025graph . These samples will be combined with large language models (LLMs) for downstream tasks such as 3D reconstruction lei2023mac ; dong2024physical ; dong2024df . In parallel, we will benchmark recent compression techniques tan2025perturbation ; liu2025rora ; li2025mutual ; tan2025harmony ; ji2025computation ; liu2025toward to support efficient deployment on embedded devices niu2025mobile ; yuan2021work ; yuan2022mobile ; ji2025computation ; liu2023scalable .
8 Related Work
LLM Agents.
@@ 8 Related Work
LLM agents unify reasoning and action. ReAct yao2023react interleaves language and actions. Toolformer schick2023toolformer self-supervises API calls, while WebGPT nakano2021webgpt reasons via web browsing. AutoGPT autogpt2023 chains subtasks autonomously. AgentBench liu2023agentbench , ReWoo xu2023rewoo , and HuggingGPT shen2023hugginggpt explore modularity and tool use. CAMEL li2023camel , ChatDev chatdev , AutoGen wu2023autogen , CrewAI crewai_github , ToolBench guo2024stabletoolbench ,and LangChain langchain_docs focus on multi-agent collaboration and LLM orchestration. GITM zhu2023ghost integrates LLMs with memory and knowledge to build generally capable agents. Although these agents show strong reasoning-action abilities, their trajectories remain challenging to compress and generalize.
@@
Distillation and Fine-Tuning. Token-level distillation hinton2015distilling ; turc2019well ; eldan2023tinystories ; gu2024minillm ; minixhofer2025cross ; cui2025multi ; zhang2024knowledge compresses LLMs via soft target alignment and teacher-guided training. Recent work improves multi-granular supervision gu2024minillm , cross-modal transfer zhou2023unidistill , and compact pretraining zhang2024tinyllama . However, these methods target static text generation, not structured reasoning-action trajectories.
@@
Trajectory Modeling and Behavioral Cloning. Recent work in sequence-level imitation wang2022self ; ke2023ccil ; swamy2022sequence ; cundy2023sequencematch ; le2016smooth ; yang2023seq2seq ; zhang2020nested , behavior cloning codevilla2018end ; hepburn2024model ; zhou2024developing ; wen2020fighting ; torabi2018behavioral ; robertson2020concurrent , and diffusion planning janner2022planning emphasizes temporal coherence in agent behavior. SpanBERT joshi2020spanbert masks contiguous spans instead of individual tokens. ConvBERT jiang2020convbert uses span-based convolutions instead of attention heads. LLM agents require structured objectives to model reasoning and action jointly—captured by our Structured Agent Distillation framework.
9 Conclusion
We propose Structured Agent Distillation, a structured compression framework that segments teacher trajectories into reasoning and action spans, enabling student agents to better mimic high-level reasoning and low-level execution beyond shallow token-level imitation.
Experiments on ALFWorld, WebShop, and HotPotQA-ReAct demonstrate consistent gains in task success, reasoning efficiency, and CoT consistency over token-level baselines. Scaling and ablation studies further highlight the value of structured supervision under limited capacity.
Our results underscore the importance of preserving agent trajectory structure for training lightweight, deployable agents, paving the way for future work in structured knowledge transfer and real-world decision-making.
@@ Appendix A Structured Agent Distillation framework
A.1 Teacher-Student Interaction Illustration
Figure 3: Illustration of the teacher-student interaction in our Structured Agent Distillation framework. The teacher model generates both the chain-of-thought (CoT) [75] reasoning steps and the corresponding action trajectory based on the task input. The student model learns to imitate these trajectories by minimizing both reasoning-level and action-level distillation losses. This joint supervision allows the student to capture both high-level problem-solving strategies and low-level task executions.
Figure 3 illustrates the core mechanism of our Structured Agent Distillation framework.
Given a task prompt, the teacher model first produces a chain-of-thought reasoning trajectory, which captures its intermediate thinking steps and decision rationale. It then generates an action trajectory, consisting of task-specific outputs (e.g., tool invocations, API calls, environment actions).
@@ A.1 Teacher-Student Interaction Illustration
A.2 Reasoning–Action Segmentation
Figure 4: Illustration of the segmented trajectory structure used in our Structured Agent Distillation framework. The teacher’s trajectory is divided into a reasoning segment (language-based, multi-step inference) and an action segment (tool invocation or final structured decision). Each segment is used to compute a dedicated imitation loss for the student model.

### v3

@@ 3.8. Curriculum Sampling in Structured Distillation
$C(\tau)=\alpha\cdot\mathrm{len}(r_{1:k})+\beta\cdot\mathrm{len}(a_{1:m})+\gamma\cdot\mathrm{entropy}\!\left(\pi_{T}(\tau)\right),$ (13)
where $\mathrm{len}(r_{1:k})$ and $\mathrm{len}(a_{1:m})$ denote the lengths of the reasoning and action segments, respectively, and $\mathrm{entropy}(\pi_{T}(\tau))$ reflects teacher uncertainty. The weights $\alpha,\beta,\gamma$ balance their relative contributions. During training, trajectories are sorted by $C(\tau)$, allowing the model to start with examples and gradually progress to more complex ones. Detailed analysis in Appendix N.
3.9. Training Algorithm
Algorithm 1 outlines our structured agent distillation process. The student policy $\pi_{\theta}$ learns to imitate the teacher $\pi_{T}$ across reasoning stages, guided by a curriculum scheduler $\mathcal{C}$ that samples increasingly complex trajectories. Each trajectory is tokenized into reasoning and action spans, and the student predicts tokens autoregressively. The objective aggregates reasoning and action losses, and gradients are backpropagated to update $\theta$.
Algorithm 1 Structured Agent Distillation (SAD)
@@ 3.9. Training Algorithm
2:for epoch =1=1 to $E$ do
3:  Sample multi-step trajectory $\tau=\{(r^{(i)},a^{(i)},o^{(i)})\}_{i=1}^{K}\sim\mathcal{C}$
4:  Linearize and tokenize $\tau^{\prime}=\texttt{[REASON]}\penalty 10000\ r^{(i)}\penalty 10000\ \texttt{[ACT]}\penalty 10000\ a^{(i)}\penalty 10000\ \texttt{[OBS]}\penalty 10000\ o^{(i)}$
5:  Construct binary masks $m_{r},m_{a},m_{o}$ over tokens $x_{1:T}$
6:  Forward pass: obtain student logits $p_{S}(\cdot\mid x_{<t})$
@@
8:  Aggregate weighted total loss $\mathcal{L}_{\text{total}}=\lambda_{r}\mathcal{L}_{\text{CoT}}+\lambda_{a}\mathcal{L}_{\text{Act}}$
9:  Update parameters $\theta\leftarrow\theta-\eta\,\nabla_{\theta}\mathcal{L}_{\text{total}}$
10:end for
4. Experiments
4.1. Experimental Setups
Agent Environments. We evaluate on three representative benchmarks: Embodied benchmark ALFWorld ALFWorld20 for embodied instruction following, Web benchmark WebShop yao2022webshop towards scalable real-world web interaction with grounded language agents, and multi-hop question benchmarks HotPotQA-ReAct yao2023react for multi-hop QA with reasoning traces. ALFWorld and WebShop test structured decision-making, while HotPotQA-ReAct requires open-ended multi-hop reasoning and free-form answer generation—thus already covering both discrete and natural-language modalities. All reported results are averaged over 5 independent runs.
Baselines. The baselines include token level KD gu2024minillm, word-level KD distilbert; lightpaff, and SeqKD (skd; alpaca; ITGPT4; lima).
Dataset Statistics. We adopt standard splits from existing ReAct-based benchmarks. ALFWorld comprises 8,055 instruction-following trajectories (5,400 train / 1,200 val / 1,475 test). WebShop contains 12,000 web interaction samples (8,000 / 2,000 / 2,000), and HotPotQA-ReAct includes 90,447 multi-hop reasoning examples (84,000 / 3,447 / 3,000).
The details of Experimental Setups in Appendix B.
4.2. Experimental Results and Analysis
@@
Table 3. Unified comparison of task success (↑), reasoning length (↓), CoT match (↑), and episode latency (steps) (↓) across ALFWorld, WebShop, and HotPotQA. The teacher is a GPT-2-1.5B ReAct-style agent. Students trained via Structured Agent Distillation consistently outperform token-level KD gu2024minillm, KD distilbert; lightpaff, and SeqKD skd; vicuna; alpaca; ITGPT4; lima baselines.
@@
Table 4. Unified comparison of task success (↑), reasoning length (↓), CoT match (↑), and episode latency (↓) across ALFWorld, WebShop, and HotPotQA. The teachers are OPT-13B, LLaMA-13B, and Orca2-13B ReAct-style agents. Students trained via Structured Agent Distillation consistently outperform token-level gu2024minillm, KD distilbert; lightpaff, and SeqKD skd; vicuna; alpaca; ITGPT4; lima baselines.
Model Task Success ↑ Reasoning Length ↓ CoT Match Rate↑ Episode Latency ↓
@@
KD (LLaMA-7B) 62.1 56.9 70.1 42.5 38.3 44.0 71.2 66.0 79.0 6.8 6.3 5.3
SeqKD (LLaMA-7B) 63.0 58.1 70.9 41.7 37.9 43.6 72.5 67.1 80.0 6.8 6.2 5.2
Token-LLaMA-7B 64.2 59.3 71.5 41.1 37.5 43.2 73.0 68.3 81.2 6.7 6.1 5.2
@@
Teacher (Orca2-13B) 78.1 75.6 84.3 37.0 34.6 39.1 100.0 100.0 100.0 6.3 5.8 4.6
KD (Orca2-7B) 64.0 59.2 72.4 41.6 37.8 43.1 73.1 68.7 81.5 6.7 6.2 5.2
SeqKD (Orca2-7B) 65.2 60.4 73.5 40.9 37.2 42.5 74.3 69.4 82.6 6.6 6.1 5.1
Token-Orca2-7B 66.3 61.7 74.8 40.3 36.9 42.0 75.6 70.2 83.4 6.6 6.0 5.1
Ours (Orca2-7B) 70.5 66.2 78.6 37.8 34.2 38.9 79.4 74.6 86.5 6.3 5.8 4.7
@@
CoT Consistency (defined in Appendix C). As shown in Table 3, our method achieves higher CoT match rates across all settings, demonstrating stronger structural alignment with teacher reasoning.
@@
Segment Mask Validation. Figure 8 (Appendix F) confirms that token-level masks align accurately with reasoning/action spans across environments.
@@
5. Scaling Analysis
Figure 3. Scaling behavior of student agents across model sizes. Top-Left: Task Success Rate (%). Top-Right: Reasoning Length (tokens). Bottom-Left: Chain-of-Thought (CoT) Match Rate (%). Bottom-Right: Episode Latency (steps). Structured Agent Distillation consistently outperforms the best baseline method (Token-level gu2024minillm) and better approaches teacher performance as model capacity increases.
[…]
Scaling behavior of student agents across model sizes.
To assess scalability, we transfer trajectories from a GPT-2-1.5B teacher model into student models with 120M, 340M, and 760M parameters using Structured Agent Distillation, and compare the outcomes against the best-performing baseline method (token-level gu2024minillm).
@@ 5. Scaling Analysis
Structured Agent Distillation scales effectively with student capacity, enhancing task success, planning efficiency, and structural reasoning alignment. Improvements are most pronounced at smaller scales (e.g., 120M, 340M), where it mitigates the performance degradation commonly observed with token-level imitation. Additional scaling results for OPT and LLaMA models are presented in Appendix G.
6. Ablation Studies
We conduct ablation studies to understand the contribution of each component in our Structured Agent Distillation framework.
6.1. Ablation on Reasoning, Action, and Segmentation Components
Specifically, we analyze the roles of reasoning supervision, action supervision, and span segmentation.
Table 5. Ablation study on ALFWorld using a 340M-parameter student model. Each variant controls whether reasoning, action supervision, and explicit span segmentation are enabled. We report task success rate, CoT match rate, and episode latency(steps).
Method Reason Action Segm Succ (%) CoT (%) Episode Latency
@@ 6.1. Ablation on Reasoning, Action, and Segmentation Components
Removing reasoning supervision ($\mathcal{L}_{\text{CoT}}$) significantly degrades CoT matches and increases latency, indicating reduced planning coherence.
@@
These results demonstrate that our method captures structurally distinct signals crucial for reasoning-action alignment.
6.2. Ablation on $\lambda_{r}$$\lambda_{a}$ Ratio
We conduct an ablation study to investigate how the balance between reasoning and action supervision affects student performance. Specifically, we vary the weighting ratio between the CoT loss ($\lambda_{\text{CoT}}$) and the action loss ($\lambda_{\text{Act}}$), while keeping the total loss weight fixed. As shown in Table 6, using both types of supervision yields the best results, with a balanced 1:1 ratio outperforming all other settings across task success rate, CoT match rate, and episode length. Interestingly, even when using only one type of supervision (either $\lambda_{\text{CoT}}$ or $\lambda_{\text{Act}}$ set to zero), our structured distillation approach still outperforms token-level baselines. These results highlight the complementary nature of reasoning and action spans in learning efficient, high-quality student agents.
Table 6. Ablation study on ALFWorld using a 340M-parameter student model. We vary the loss weight ratios between reasoning (CoT) and action supervision while keeping their sum fixed. Structured Agent Distillation consistently outperforms token-level baselines across all settings.
CoT : Act Ratio Succ (%) CoT (%) Episode Latency
1.0 : 0.0 53.4 69.2 8.2
0.0 : 1.0 52.7 66.1 8.3
0.5 : 1.0 55.8 70.3 7.7
1.0 : 1.0 56.3 71.5 7.1
2.0 : 1.0 56.1 70.8 7.3
Token-Level Baseline 52.1 68.1 8.6
Ablation study on curriculum sampling in Appendix N, Qualitative Examples and Faithfulness in Appendix O.
7. Discussion
@@
Loss Design and Supervision. Although the total loss combines $\mathcal{L}_{\text{CoT}}$ and $\mathcal{L}_{\text{Act}}$, ablation studies show that removing either term causes clear degradation in CoT alignment, task success, and planning efficiency. This demonstrates that span-level supervision offers complementary signals beyond token-level KL objectives.
@@
[…]
8. Related Work
@@
LLM Agents. LLM agents unify reasoning and action. ReAct yao2023react interleaves language and actions. Toolformer schick2023toolformer self-supervises API calls, while WebGPT nakano2021webgpt reasons via web browsing. AutoGPT autogpt2023 chains subtasks autonomously. AgentBench liu2023agentbench, ReWoo xu2023rewoo, and HuggingGPT shen2023hugginggpt explore modularity and tool use. CAMEL li2023camel, ChatDev chatdev, AutoGen wu2023autogen, CrewAI crewai_github, ToolBench guo2024stabletoolbench, and LangChain langchain_docs focus on multi-agent collaboration and LLM orchestration. GITM zhu2023ghost integrates LLMs with memory and knowledge to build generally capable agents. Although these agents show strong reasoning-action abilities, their trajectories remain challenging to compress and generalize.
@@
Distillation and Fine-Tuning. Token-level distillation hinton2015distilling; eldan2023tinystories; gu2024minillm; minixhofer2025cross; cui2025multi; zhang2024knowledge compresses LLMs via soft target alignment and teacher-guided training. Recent work improves multi-granular supervision gu2024minillm, cross-modal transfer zhou2023unidistill, and compact pretraining zhang2024tinyllama. However, these methods target static text generation, not structured reasoning-action trajectories.
@@
Trajectory Modeling and Behavioral Cloning. Recent work in sequence-level imitation wang2022self; ke2023ccil; swamy2022sequence; cundy2023sequencematch; le2016smooth; yang2023seq2seq; zhang2020nested, behavior cloning codevilla2018end; hepburn2024model; zhou2024developing; wen2020fighting; torabi2018behavioral; robertson2020concurrent emphasizes temporal coherence in agent behavior. SpanBERT joshi2020spanbert masks contiguous spans instead of individual tokens. ConvBERT jiang2020convbert uses span-based convolutions instead of attention heads. LLM agents require structured objectives to model reasoning and action jointly—captured by our Structured Agent Distillation framework.
9. Conclusion
We present Structured Agent Distillation, a compression framework that segments teacher trajectories into reasoning and action spans, allowing students to better replicate high-level reasoning and low-level execution beyond token-level imitation. Our method achieves consistent gains in task success, reasoning efficiency, and CoT alignment over token-level baselines. Scaling and ablation results confirm the benefits of structured supervision under limited capacity. These findings highlight the importance of preserving trajectory structure for training lightweight agents and open avenues for structured knowledge transfer in real-world decision making.
@@ Appendix A Structured Agent Distillation framework
A.1. Teacher-Student Interaction Illustration
Figure 4. Illustration of the teacher-student interaction in our Structured Agent Distillation framework. The teacher model generates both the chain-of-thought (CoT) ott2023thoughtsource reasoning steps and the corresponding action trajectory based on the task input. The student model learns to imitate these trajectories by minimizing both reasoning-level and action-level distillation losses. This joint supervision allows the student to capture both high-level problem-solving strategies and low-level task executions.
[…]
An overview of the loss flow in our Structured Agent Distillation framework. The student model learns from the teacher’s trajectory by decomposing the learning signals into reasoning and action objectives. These loss terms are aggregated to guide effective student optimization.
Figure 4 illustrates the core mechanism of our Structured Agent Distillation framework. Given a task prompt, the teacher model first produces a chain-of-thought reasoning trajectory, which captures its intermediate thinking steps and decision rationale. It then generates an action trajectory, consisting of task-specific outputs (e.g., tool invocations, API calls, environment actions).
@@ A.1. Teacher-Student Interaction Illustration
A.2. Reasoning–Action Segmentation
Figure 5. Illustration of the segmented trajectory structure used in our Structured Agent Distillation framework. The teacher’s trajectory is divided into a reasoning segment (language-based, multi-step inference) and an action segment (tool invocation or final structured decision). Each segment is used to compute a dedicated imitation loss for the student model.
[…]
Illustration of the segmented trajectory structure used in our Structured Agent Distillation framework. The teacher’s trajectory is divided into a reasoning segment (language-based, multi-step inference) and an action segment (tool invocation or final structured decision). Each segment is used to compute a dedicated imitation loss for the student model.